# On the Invariant and Geometric Structure of the Holomorphic Unified Field Theory

### v1

@@ Abstract
We present the invariant structure of a Holomorphic Unified Field Theory in which gravity and gauge interactions arise from a single geometric framework. The theory is formulated on a product principal bundle, with one connection, and curvature equipped with a Hermitian field on a complexification of spacetime. From a single $\mathrm{Diff}(M)\times G$–invariant action, variation yields the Einstein and Yang–Mills equations together with their paired Bianchi identities. A compatibility condition is implemented either definitionally or through an auxiliary penalty functional. It enforces that the antisymmetric part of our Hermitian field is exactly the gauge field’s curvature on the real slice.
My methods are really methods of working and thinking; this is why they have crept in everywhere anonymously.
Emmy Noether, Letter to Helmut Hasse (1931)
@@ 1 Introduction
In 1945, Albert Einstein wrote, “Every attempt to establish a unified-field theory must start, in my opinion, from a group of transformations.” This means that unification is a symmetry-first program [1]. Emmy Noether had the view that any candidate must be built from a single invariant action [2]. From that one symmetry comes conserved currents and differential identities and or Bianchi-type relations [2, 3]. Noether used general relativity (GR) as the archetype for an infinite-dimensional symmetry diffeomorphisms, showing how conservation becomes a consequence of symmetry, not an add-on. The Holomorphic Unified Field Theory (HUFT) implements Einstein’s symmetry-first mandate by choosing one transformation structure and building geometry via a single metric-compatible connection plus a single gauge connection. It satisfies Noether by using one invariant action whose variations give both the Einstein and Yang–Mills equations. At the same time, the same symmetry yields Bianchi identities and conserved currents [4, 5, 6, 7]. HUFT solves the dynamical part of unification in its own sense with a single UV connection and coupling, regulated variational mechanism [8, 9, 10]. It does not yet solve the unique GUT-style unification unless we add and prove a geometric or topological locking principle that fixes the relative normalization and uniquely determines the internal group [11, 12].
@@
We develop a symmetry–first formulation of HUFT, whose aim is to package gravity and gauge interactions into a single geometric structure without introducing new propagating degrees of freedom. The guiding principle is that dynamics should follow from one invariant action built from one set of geometric data on a single underlying bundle, so that the familiar Einstein and Yang–Mills sectors are recovered as two aspects of the same symmetry [5, 6]. In this sense, unification here means shared kinematics, identities, and variational origin is not a proliferation of fields or ad hoc couplings.
@@
We work with a product principal bundle carrying a single connection that contains both the spin connection and the internal gauge connection [5]. We also introduce a Hermitian packaging of the spacetime metric with an antisymmetric two–form valued in the gauge algebra. Crucially, that antisymmetric part is not given independent dynamics, either by definition or through an algebraic compatibility term in the action, it is tied pointwise to the gauge curvature on the real slice of the theory [7]. All index operations and geometric constructions are performed with the ordinary Lorentzian metric [3]. This ensures that the holomorphic language functions as a bookkeeping device that makes the unification transparent while remaining mathematically safe.
@@
We obtain from this setup two core results; first, at the classical level, the unified theory on the real slice is equivalent to the standard system of general relativity plus Yang–Mills with the same matter content [3, 4]. The conservation laws and Bianchi identities arise as a single Noether identity associated with diffeomorphism and internal gauge invariance [2, 7]. Second, at the level of formal path integrals, integrating out the auxiliary two–form reduces the functional measure to that of the standard theory, indicating that no additional propagating modes or counterterms are introduced by the holomorphic packaging [10, 13, 14]. We are explicit about the status of this second statement, as it follows established physics practice but is not presented as a measure–theoretic proof.
@@
Beyond demonstrating equivalence, the paper advances two structural claims. The first is a minimality statement that if one insists on diffeomorphism– and gauge–natural, second–order dynamics built from the unified geometric data, the Einstein–Hilbert and Yang–Mills terms are singled out up to boundary and topological contributions [3, 6]. The second is an internal–group constraint, the holomorphic and Yang–Mills structure, together with chirality and charge integrality, points toward a unitary internal group with the minimal viable choice breaking to the Standard Model on the real slice [11, 12, 15].
@@
While the mathematical structures we employ, Hermitian frame bundles, spin bundles, and holomorphic stability criteria, are well established in differential and complex geometry [16, 17], the novelty of this work lies in their assembly into a single unification framework. The complexified metric $g=h+iB$ is interpreted such that its antisymmetric part is constrained to equal the Yang-Mills curvature, thereby formulating gravitational and gauge data into one Hermitian structure. This identification, together with the formulation on the product bundle $P_{Spin(1,3)}\times_{M}P_{G}$ yields a single invariant action reproducing Einstein-Hilbert and Yang-Mills dynamics. The originality of the theory is not in introducing new mathematical objects, but in recombining standard geometric ingredients into a symmetry-driven unification principle. The Hermitian metric is constrained pointwise to be the Yang-MMills curvature, so that gravity and gauge interactions are implemented in a single Hermitian structure, while remaining dynamically equivalent to Einstein-Hilbert and plus Yang-Millss. In this sense, the work is in spirit symmetry based on Noetherian unification principles, employing a holomorphic bundle formalism.
@@
Our goal is therefore twofold, to give a self–contained account of the invariant structure that makes the unification precise and mathematically consistent, and to delineate the exact sense in which the construction is equivalent to known physics while clarifying what is genuinely new and what is not. The remainder of the paper spells out the geometric data and symmetries, derives the field equations and identities from a single action, proves the equivalence at the classical level, analyses the formal quantum reduction, and discusses the minimality and internal–group considerations, together with open directions on anomaly data, matter–bundle building, and measure–theoretic rigour. We believe that this can and should be compared with past theories to induce a discussion about how a unified theory should be interpreted and formulated [18, 19, 20, 21, 22, 23, 24, 25].
@@ 2 Invariant Geometry of HUFT
We will show that the Hermitian metric as given by HUFT [26, 27, 28] does satisfy a unified theory as described by Albert Einstein [1] and Emmy Noether [2]. We let $M$ be an oriented, time–orientable, spin 44–manifold on the real slice. We let $P_{\mathrm{Spin}}\to M$ be the $\mathrm{Spin}(1,3)$ frame bundle and $P_{G}\to M$ a principal GG–bundle. We define the product principal bundle:
$P_{\mathrm{tot}}\;=\;P_{\mathrm{Spin}}\times_{M}P_{G},\qquad H:=\mathrm{Spin}(1,3)\times G,\qquad\mathfrak{h}=\mathfrak{spin}(1,3)\oplus\mathfrak{g}.$ (1)
Now we let $h\in\Gamma(S^{2}T^{*}M)$ be the smooth sections of the bundle of symmetric 2-forms with signature (+++)(-+++) be a Lorentzian metric on $M$ and $\mathcal{A}=(\omega,A)$ an HH–connection on $P_{\mathrm{tot}}$, with curvatures:
@@
where $R\in\Omega^{2}(M,\mathfrak{spin}(1,3))$ is the Riemann curvature 22–form of $\omega$ and $F\in\Omega^{2}(M,\mathrm{ad}P_{G})$ is the Yang–Mills field strength of $A$.
@@
We now derive the internal group from the fiber geometry. We have $E\to M$ as the matter bundle of rank $n$ with Hermitian metric $h_{E}$, the first Chern class of the complex vector bundle is zero $c_{1}(E)=0$, the determinant line bundle $\text{det}E$ is topologically trivial so no net U(1) flux meaning the structure group can drop from $U(n)$ to $SU(n)$, a nowhere-vanishing holomorphic volume form $\Omega_{E}\in H^{0}(M,\wedge^{n}E^{*})$ meaning that there exists a global holomorphic section of the determinant line bundle $\wedge^{n}E^{*}\cong(\det E)^{*}$ which never vanishes on $M$ ion local holomorphic frames $\{e_{i}\}_{i=1}^{n}$ of $E$, the form can be written as $\Omega_{E}=f(z)\,e^{1}\wedge\cdots\wedge e^{n}$ with a holomorphic function $f(z)\neq 0$ everywhere. The existence of such a nowhere–vanishing section trivializes the determinant line bundle, implying $\det E\simeq\mathcal{O}_{M}$ and hence $c_{1}(E)=0$. Geometrically, $\Omega_{E}$ plays the role of a holomorphic volume element on the fibers of $E$, fixing an intrinsic orientation and allowing us to identify totally antisymmetric tensors via the holomorphic epsilon symbol. The subgroup of $GL(n,\mathbb{C})$ preserving both the Hermitian metric $h_{E}$ and $\Omega_{E}$ is $SU(n)$ and thus the data $(E,h_{E},\Omega_{E})$ define an $SU(n)$-structure on the bundle. Physically, this reduction eliminates the overall $U(1)$ determinant factor, enforcing traceless gauge transformations and ensuring charge quantization in the $SU(n)$ fiber, as required for unified models such as $SU(5)$. And we assume slope-stable so the compatible connection is Hermitian–Yang–Mills (HYM) [40, 41]. Preservation of $(h_{E},\Omega_{E})$ reduces the structure group of $E$ to $\mathrm{SU}(n)$, and the HYM connection has holonomy in $\mathrm{SU}(n)$. We let $Y\in\mathfrak{su}(n)$ be the traceless character compatible with $\Omega_{E}$. We normalize $Y$ by the calibrated pairing so that:
@@
fixing the abelian charge units without convention. Imposing a single adjoint breaking on the real slice, the hypercharge integrality above, and chiral, anomaly-free matter for one family from associated bundles of $E$, the minimal rank is $n=5$ [12, 15]. Therefore, the maximal compatible internal symmetry is:
@@
under a holomorphic adjoint ($\mathbf{24}$) reduction, with the standard hypercharge generator [11, 12]. In what follows, we take $G=\mathrm{SU}(5)$ above the mass-energy scale $M_{*}$, with a holomorphic adjoint reduction to the Standard Model group on the real slice. Given the rank-$n$ holomorphic bundle $E\to M$ with Hermitian metric on the internal vector bundle $h_{E}$ and a nowhere-vanishing holomorphic volume form $\Omega_{E}\in H^{0}(M,\wedge^{n}E^{*})$, the structure group reduces as:
@@
The preservation of $h_{E}$ picks unitary changes of frame, such as $\mathrm{U}(n)\subset\mathrm{GL}(n,\mathbb{C})$. Preservation of $\Omega_{E}$ forces $\det=1$ on $\mathrm{U}(n)$, giving $\mathrm{SU}(n)$. If, in addition, $c_{1}(E)=0$ and $E$ is slope-stable, the unique compatible connection is Hermitian–Yang–Mills and its holonomy lies in $\mathrm{SU}(n)$ [40, 41]. The internal gauge symmetry is the automorphism group preserving $(h_{E},\Omega_{E})$, namely $G=\mathrm{SU}(n)$. Imposing chirality and anomaly cancellation for one SM family, hypercharge integrality via $\frac{1}{8\pi^{2}}\int\mathrm{Tr}(F_{Y}\wedge F_{Y})\in\mathbb{Z}$, and a single adjoint breaking on the real slice fixes the minimal rank to $n=5$, we have:

### v2

# Holomorphic Unified Field Theory; Hermitian metric; Einstein–Yang–Mills; diffeomorphism invariance; gauge invariance; product principal bundle; Bianchi identities; curvature compatibility

@@ Abstract
We present the invariant structure of a Holomorphic Unified Field Theory in which gravity and gauge interactions arise from a single geometric framework. The theory is formulated using a product principal bundle, with one connection, and curvature equipped with a Hermitian field on a complexification of spacetime. From a single $\mathrm{Diff}(M)\times G$-invariant action, variation yields the Einstein and Yang–Mills equations together with their paired Bianchi identities. A compatibility condition is implemented either definitionally or through an auxiliary penalty functional. It enforces that the antisymmetric part of our Hermitian field is the gauge field’s exact curvature on the real slice.
keywords:
Holomorphic Unified Field Theory; Hermitian metric; Einstein–Yang–Mills; diffeomorphism invariance; gauge invariance; product principal bundle; Bianchi identities; curvature compatibility
[…]
@@ 1 Introduction
In 1945, Albert Einstein wrote, “Every attempt to establish a unified-field theory must start, in my opinion, from a group of transformations.” This means that unification is a symmetry-first program Einstein1945 . Emmy Noether had the view that any candidate must be built from a single invariant action NoetherTrans . From that one symmetry, conserved currents and differential identities and or Bianchi-type relations originate NoetherTrans ; WaldGR . Noether used general relativity (GR) as the archetype for infinite-dimensional symmetry diffeomorphisms, showing how conservation becomes a consequence of symmetry, not an add-on. The Holomorphic Unified Field Theory (HUFT) implements Einstein’s symmetry-first mandate by choosing one transformation structure and building geometry via a single metric-compatible connection plus a single gauge connection. It satisfies Noether by using one invariant action whose variations give both the Einstein and Yang–Mills equations. At the same time, the same symmetry yields Bianchi identities and conserved currents YangMills1954 ; KobayashiNomizu1 ; Nakahara ; Trautman1979 . HUFT solves the dynamical part of unification in its own sense with a single UV connection and coupling, regulated variational mechanism Slavnov1972 ; Taylor1971 ; BRST1976 . It does not yet solve the unique GUT-style unification unless we add and prove a geometric or topological locking principle that fixes the relative normalization and uniquely determines the internal group GeorgiGlashow1974 ; Slansky1981 .
In this paper, we make the following claims precise, that a symmetry-first formulation on a product principal bundle with a single connection reproduces Einstein–Hilbert and Yang–Mills dynamics from one invariant action, together with the paired Noether/Bianchi identities. On the real slice, the Hermitian packaging $g=h+iB$ is a kinematic unification device whose antisymmetric sector is constrained to coincide with the gauge curvature, $B\equiv F$, and introduces no additional propagating degrees of freedom. We prove classical equivalence solution-space bijection modulo $\mathrm{Diff}(M)\times G$ between the unified real-slice system and standard Einstein–Yang–Mills (EYM) and matter. We state the minimality/uniqueness claim for diffeo- and gauge-natural second-order couplings, up to boundary or topological terms, and delineate what remains open, a truly unique internal-group selection without an additional locking principle.
@@
We develop a symmetry-first formulation of HUFT, whose aim is to package gravity and gauge interactions into a single geometric structure without introducing new propagating degrees of freedom. The guiding principle is that dynamics should follow from one invariant action built from one set of geometric data on a single underlying bundle, so that the familiar Einstein and Yang–Mills sectors are recovered as two aspects of the same symmetry KobayashiNomizu1 ; Nakahara . In this sense, unification here means shared kinematics, identities, and variational origin is not a proliferation of fields or ad hoc couplings.
@@
We work with a product principal bundle carrying a single connection that contains both the spin connection and the internal gauge connection KobayashiNomizu1 . We also introduce a Hermitian packaging of the spacetime metric with an antisymmetric two–form valued in the gauge algebra. Crucially, that antisymmetric part is not given independent dynamics, either by definition or through an algebraic compatibility term in the action; it is tied pointwise to the gauge curvature on the real slice of the theory Trautman1979 . All index operations and geometric constructions are performed with the ordinary Lorentzian metric WaldGR . This ensures that the holomorphic language functions as a bookkeeping device that makes the unification transparent while remaining mathematically safe.
@@
We obtain from this setup two core results first, at the classical level, the unified theory on the real slice is equivalent to the standard system of general relativity plus Yang–Mills with the same matter content WaldGR ; YangMills1954 . The conservation laws and Bianchi identities arise as a single Noether identity associated with diffeomorphism and internal gauge invariance NoetherTrans ; Trautman1979 . Second, at the level of formal path integrals, integrating out the auxiliary two–form reduces the functional measure to that of the standard theory, indicating that no additional propagating modes or counterterms are introduced by the holomorphic packaging BRST1976 ; DeWitt1965 ; Vassilevich2003 . We are explicit about the status of this second statement, as it follows established physics practice but is not presented as a measure–theoretic proof.
@@
Beyond demonstrating equivalence, this paper advances two structural claims. The first is a minimality statement that if one insists on diffeomorphism- and gauge-natural, second-order dynamics built from the unified geometric data, the Einstein–Hilbert and Yang–Mills terms are singled out up to boundary and topological contributions WaldGR ; Nakahara . The second is an internal-group constraint; the holomorphic and Yang–Mills structure, together with chirality and charge integrality, points toward a unitary internal group with the minimal viable choice breaking to the Standard Model on the real slice GeorgiGlashow1974 ; Slansky1981 ; Bardeen1969 .
@@
While the mathematical structures we employ, Hermitian frame bundles, spin bundles, and holomorphic stability criteria, are well established in differential and complex geometry, the novelty of this work lies in their assembly into a single unification framework. The complexified metric $g=h+iB$ is interpreted such that its antisymmetric part is constrained to equal the Yang–Mills curvature, thereby formulating gravitational and gauge data into one Hermitian structure. This identification, together with the formulation on the product bundle $P_{Spin(1,3)}\times_{M}P_{G}$, yields a single invariant action reproducing Einstein–Hilbert and Yang–Mills dynamics. The originality of the theory is not in introducing new mathematical objects, but in recombining standard geometric ingredients into a symmetry-driven unification principle. The Hermitian metric is constrained pointwise to be the Yang–Mills curvature, so that gravity and gauge interactions are implemented in a single Hermitian structure, while remaining dynamically equivalent to Einstein–Hilbert and plus Yang–Mills. In this sense, the work is, in spirit symmetry, based on Noetherian unification principles, employing a holomorphic bundle formalism.
@@
Our goal is therefore twofold—to give a self–contained account of the invariant structure that makes the unification precise and mathematically consistent, and to delineate the exact sense in which the construction is equivalent to known physics while clarifying what is genuinely new and what is not. The remainder of the paper spells out the geometric data and symmetries, derives the field equations and identities from a single action, proves the equivalence at the classical level, analyses the formal quantum reduction, and discusses the minimality and internal-group considerations, together with open directions on anomaly data, matter-bundle building, and measure-theoretic rigour. We believe that this can and should be compared with past theories to induce a discussion about how a unified theory should be interpreted and formulated SusskindHologram1995 ; SusskindLandscape2003 ; Witten1995 ; WittenAdS1998 ; PenroseTwistor1976 ; KakuKikkawaTrees1974 ; KakuKikkawaLoops1974 ; KakuClosedSFT1988 .
@@ 2 Invariant Geometry of HUFT
We will show that the Hermitian metric as given by HUFT MoffatThompsonHUFT2025 ; MoffatThompsonFiniteNonlocal2025 ; MoffatThompsonMassSpectrum2025 ; ThompsonTopQuark2025 ; MT:ReplyToCline ; MT:GI2025 ; MT:AdSdS2025 ; T:DarkMatter2025 does satisfy a unified theory as described by Albert Einstein Einstein1945 and Emmy Noether NoetherTrans . We let $M$ be an oriented, time-orientable, spin 44–manifold on the real slice. We let $P_{\mathrm{Spin}}\to M$ be the $\mathrm{Spin}(1,3)$ frame bundle and $P_{G}\to M$ a principal $G$-bundle. We define the product principal bundle:
$P_{\rm tot}=P_{\rm Spin}\times_{M}P_{G},\quad H:=\text{Spin}(1,3)\times G,\quad\mathfrak{h}:=\mathfrak{spin}(1,3)\oplus\mathfrak{g}.$ (1)
Now, we let $h\in\Gamma(S^{2}T^{*}M)$ be a smooth section of the bundle of symmetric two-forms with signature (+++)(-+++) being a Lorentzian metric on $M$ and $\mathcal{A}=(\omega,A)$ an HH–connection on $P_{\mathrm{tot}}$, with the following curvatures:
@@
where $R\in\Omega^{2}(M,\mathfrak{spin}(1,3))$ is the Riemann curvature $2$-form of $\omega$ and $F\in\Omega^{2}(M,\mathrm{ad}P_{G})$ is the Yang–Mills field strength of $A$. We separate the bundle objects used in HUFT into internal or matter bundle data on the real slice and a holomorphic thickening used as a bookkeeping device. We take a complex rank-$n$ vector bundle $E\to M$ equipped with a holomorphic structure equivalent to a $\bar{\partial}$-operator on $E$, a Hermitian fiber metric $h_{E}$, and a nowhere-vanishing holomorphic volume form $\Omega_{E}\in H^{0}(M,\wedge^{n}E^{\ast})$. The pair $(h_{E},\Omega_{E})$ reduces the structure group from $GL(n,\mathbb{C})$ to $SU(n)$, and the associated unitary frame bundle defines a principal $SU(n)$-bundle $P\to M$ with a compatible connection $A$. When needed, we assume a complexification $\iota:M\hookrightarrow M_{\mathbb{C}}$ and an extension of the bundle to a holomorphic bundle $E^{1,0}\to M_{\mathbb{C}}$ restricting to $E$ on $M$. This thickening is used only to organize fields and does not alter topological constraints such as $c_{1}(E)=0$ imposed on the real slice. To review the first Chern class $c_{1}(E)\in H^{2}(M,\mathbb{Z})$ is defined by $c_{1}(E):=c_{1}(\det E)$. Equivalently, for any unitary connection $A$ on $E$ with curvature $F_{A}$, the Chern–Weil form $(i/2\pi)\text{Tr}(F_{A})$ is closed and represents $c_{1}(E)$ in de Rham cohomology. In particular, $\det E\simeq\mathcal{O}_{M}$, such as for $SU(r)$ bundles, implies $c_{1}(E)=0$. The class $c_{1}(E)\in H^{2}(M,\mathbb{Z})$ measures the net $U(1)$ twisting of the complex vector bundle $E$. Equivalently, it measures the twisting of the determinant line bundle $\det E:=\wedge^{\mathrm{rk}(E)}E$, since $c_{1}(E):=c_{1}(\det E)$. A useful gauge-theoretic intuition is that $c_{1}(E)$ is the quantized trace curvature or flux; as for any unitary connection $A$ on $E$ with curvature $F_{A}$, the closed $2$-form $\frac{i}{2\pi}\text{Tr}(F_{A})$ represents $c_{1}(E)$ in de Rham cohomology, and its integral over any closed $2$-cycle is an integer. In particular, $\det E\simeq\mathcal{O}_{M}$, such as for the $SU(r)$ bundles, implies $c_{1}(E)=0$, meaning that there is no net trace ($U(1)$) twisting. Here, $E^{1,0}\to M_{\mathbb{C}}$ denotes a holomorphic extension thickening of $E\to M$ to the complexified manifold, and $\mathcal{O}_{M}$ and $\mathcal{O}_{M_{\mathbb{C}}}$ denote the trivial respective holomorphic line bundles on $M$ and $M_{\mathbb{C}}$. We write $\det E:=\wedge^{n}E$ to represent the determinant line bundle. Using $M_{\mathbb{C}}$, we introduce a complex Hermitian two-tensor:
$g_{\mu\nu}:=h_{\mu\nu}+i\ell_{\ast}^{\,2}B_{\mu\nu},$ (3)
where $h_{\mu\nu}$ is the Lorentzian metric on $M$ and $B\in\Omega^{2}(M,\mathrm{ad}\,P_{G})$ is an adjoint-valued two-form pulled back from the real slice. A compatibility mechanism enforces $B=F$ on $M$, so this Hermitian field is a kinematic packaging of $(h,F)$ rather than an additional propagating sector.
In ordinary general relativity, the metric $h$ determines the Levi–Civita connection $\nabla^{\rm LC}(h)$ and, hence, the Riemann curvature two-form $R(h)$ by differentiation, schematically $h\mapsto\Gamma(h)\mapsto R(h)$. In HUFT, the unification has a different structural flavor, as the internal curvature two-form is already present at the level of the generalized Hermitian metric variable. On the real slice, the compatibility mechanism yields $B=F$, so the antisymmetric sector of the Hermitian packaging field coincides with the Yang–Mills curvature two-form. In this sense, part of the curvature data is encoded in the generalized metric variable rather than being derived only from it, and the metric/connection/curvature separation familiar from GR is reorganized.
@@
We now derive the internal group from the fiber geometry. We have $E\to M$ as the matter bundle of the rank $n$ holomorphic bundle over spacetime $M$ with Hermitian metric $h_{E}$. The first Chern class of the complex vector bundle is zero $c_{1}(E)=0$, the determinant line bundle $\text{det}E$ is topologically trivial with no net U(1) flux, meaning the structure group can drop from $U(n)$ to $SU(n)$, a nowhere-vanishing holomorphic volume form $\Omega_{E}\in H^{0}(M,\wedge^{n}E^{*})$, meaning that there exists a global holomorphic section of the determinant line bundle $\wedge^{n}E^{*}\cong(\det E)^{*}$ which never vanishes on $M$ in local holomorphic frames $\{e_{i}\}_{i=1}^{n}$ of $E$, the form can be written as $\Omega_{E}=f(z)\,e^{1}\wedge\cdots\wedge e^{n}$ with a holomorphic function $f(z)\neq 0$ everywhere.
We let $P\to M$ be a principal U(n)-bundle encoding the internal gauge symmetry $G\subseteq U(n)$. For a faithful unitary representation $\rho:G\hookrightarrow U(n)$, define the associated rank-n complex vector bundle:
$E\;:=\;P\times_{\rho}\mathbb{C}^{n}\;\xrightarrow{\;\pi\;}\;M,$ (4)
equipped with a Hermitian fiber metric $h_{E}$. The first Chern class:
$c_{1}(E)\;=\;c_{1}(\det E)\in H^{2}(M,\mathbb{Z})$ (5)
is a topological invariant of the matter bundle $E$ not of the tangent bundle, we impose
$c_{1}(E)=0,$ (6)
which is equivalent to $\det(E)$ being topologically trivial and hence reduces the structure group from $U(n)$ to $SU(n)$, so a compatible unitary connection satisfies $\mathrm{Tr}\,F_{A}=0$. This is the precise sense in which the fiber geometry selects an internal $SU(n)$ structure. We use a holomorphic thickening $M\hookrightarrow M_{\mathbb{C}}$ to organize the field content and action, but all global topological constraints, such as $c_{1}(E)=0$, are imposed on the real slice M. When needed, we assume $E$ extends to a holomorphic bundle $E^{1,0}\to M_{\mathbb{C}}$ with $\det(E^{1,0})\simeq\mathcal{O}_{M_{\mathbb{C}}}$, whose restriction to $M$ is $E$. The thickening does not weaken the global condition $c_{1}(E)=0$ on $M$. The existence of such a nowhere-vanishing section trivializes the determinant line bundle, implying $\det E\simeq\mathcal{O}_{M}$ with $\mathcal{O}_{M}$ is the trivial line object, and hence $c_{1}(E)=0$, $\det E$ is the determinant line bundle of the vector bundle. Geometrically, $\Omega_{E}$ plays the role of a holomorphic volume element on the fibers of $E$, fixing an intrinsic orientation and allowing us to identify totally antisymmetric tensors via the holomorphic epsilon symbol. The subgroup of $GL(n,\mathbb{C})$ preserving both the Hermitian metrics $h_{E}$ and $\Omega_{E}$, is $SU(n)$ and thus the data $(E,h_{E},\Omega_{E})$ define an $SU(n)$-structure on the bundle. Physically, this reduction eliminates the overall $U(1)$ determinant factor, enforcing traceless gauge transformations and ensuring charge quantization in the $SU(n)$ fiber, as required for unified models such as $SU(5)$, and we assume that this is slope-stable. Let $(M,\omega)$ be a compact Kähler manifold and $E\to M$ be a holomorphic vector bundle. The $\omega$–degree and slope are:
$\deg_{\omega}(E)=\int_{M}c_{1}(E)\wedge\frac{\omega^{m-1}}{(m-1)!},\qquad\mu_{\omega}(E)=\frac{\deg_{\omega}(E)}{\mathrm{rk}\,E}.$ (7)
We say $E$ is Mumford–Takemoto slope-stable if, for every coherent subsheaf, $0\neq F\subsetneq E$ with $0<\mathrm{rk}\,F<\mathrm{rk}\,E$ one has $\mu_{\omega}(F)<\mu_{\omega}(E)$; semistable if $\leq$; and polystable if it is a direct sum of stable bundles of the same slope. When $c_{1}(E)=0$, we have $\mu_{\omega}(E)=0$ for all $\omega$. By the Donaldson–Uhlenbeck–Yau theorem, $E$ admits a Hermitian–Yang–Mills connection if $E$ is polystable; in the $c_{1}(E)=0$ case, this means $i\,\Lambda_{\omega}F_{A}=0$ and $\mathrm{Tr}\,F_{A}=0$, so the structure group reduces to $SU(n)$. We should note that Donaldson–Uhlenbeck–Yau is a theorem about holomorphic bundles on a compact Kähler base. Accordingly, when we invoke slope-(poly)stability and the existence of a Hermitian–Yang–Mills (HYM) connection, we apply the theorem on an auxiliary compact Kähler manifold $(X,\omega)$ carrying the relevant holomorphic bundle data, such as a compact Euclideanized slice and or a compactification used for holomorphic classification. For a holomorphic bundle $E\to X$, the $\omega$-degree and slope are $\deg_{\omega}(E)=\int_{X}c_{1}(E)\wedge\omega^{m-1}/(m-1)!$ and $\mu_{\omega}(E)=\deg_{\omega}(E)/\mathrm{rk}\,E$. If $E$ is polystable, DUY implies the existence of an HYM connection; when $c_{1}(E)=0$, this includes $\mathrm{Tr}\,F_{A}=0$ and $i\Lambda_{\omega}F_{A}=0$. The physical spacetime in our field theory remains the Lorentzian real slice $(M,h)$, on which no Kähler hypothesis is imposed; the DUY input is used only as a standard existence or uniqueness criterion for the compatible unitary connection associated with the holomorphic internal or matter bundle data. so the compatible connection is Hermitian–Yang–Mills (HYM) Donaldson1985 ; UhlenbeckYau1986 . The classical Donaldson–Uhlenbeck–Yau theorem is stated for compact Kähler manifolds. For our purposes, it suffices to use the Hermitian non-Kähler generalizations of the Kobayashi–Hitchin correspondence. We let $(X,\omega)$ be a compact complex manifold equipped with a Gauduchon metric, such as $\partial\bar{\partial}\,\omega^{n-1}=0$. Then, a holomorphic vector bundle $E\to X$ is polystable with its slope defined using $\omega$ if and only if $E$ admits a Hermitian–Einstein equivalently Hermitian–Yang–Mills connection $A$ satisfying:
$F_{A}^{0,2}=0,\qquad\sqrt{-1}\,\Lambda_{\omega}F_{A}=\lambda_{\mathrm{HE}}\,\mathrm{Id}_{E},$ (8)
Here, $\Lambda_{\omega}$ is the contraction adjoint to wedging with $\omega$ equivalently, for a $(1,1)$-form $\alpha$, $\Lambda_{\omega}\alpha=\omega^{i\bar{j}}\alpha_{i\bar{j}}$ in local holomorphic coordinates, and $F_{A}^{0,2}=0$ means $A$ defines a holomorphic structure on $E$, and $\lambda_{\mathrm{HE}}\in\mathbb{R}$ is the Hermitian–Einstein constant determined by the slope of $E$:
$\lambda_{\mathrm{HE}}=2\pi\,\mu_{\omega}(E)/\mathrm{Vol}_{\omega}(M)$ (9)
with $\mu_{\omega}(E)=\deg_{\omega}(E)/\mathrm{rk}(E)$ and:
$\deg_{\omega}(E)=\int_{M}c_{1}(E)\wedge\omega^{n-1}/(n-1)!.$ (10)
In particular, if $\det E\simeq\mathcal{O}_{M}$ so $c_{1}(E)=0$, then $\lambda_{\mathrm{HE}}=0$.
In the $SU(n)$ case in particular, when $c_{1}(E)=0$, one has $\lambda=0$. Thus, the existence or uniqueness input we use does not require the Kähler condition, but only a compact Hermitian Gauduchon background. Preservation of $(h_{E},\Omega_{E})$ reduces the structure group of $E$ to $\mathrm{SU}(n)$, and the HYM connection has holonomy in $\mathrm{SU}(n)$. We let $Y\in\mathfrak{su}(n)$ be the traceless character compatible with $\Omega_{E}$. We normalize $Y$ by the calibrated pairing, to calibrate the pairing here we pick the invariant Lie-algebra inner product, the trace or Killing form, and fix the normalization of the $U(1)$ generator $Y$ so that the Chern–Weil pairing with the curvature lands in the integer lattice, so that:
@@
fixing the abelian charge units without convention. Imposing a single adjoint breaking on the real slice, the hypercharge integrality above, and chiral, anomaly-free matter for one family from associated bundles of $E$, the minimal rank is $n=5$ Slansky1981 ; Bardeen1969 . A simple necessary condition comes from the Lie-group rank. The Standard Model gauge group has
$\text{rank}\bigl(SU(3)_{c}\times SU(2)_{L}\times U(1)_{Y}\bigr)$ (12)
$=\text{rank}(SU(3))+\text{rank}(SU(2))+\text{rank}(U(1))=2+1+1=4.$ (13)
For any connected compact Lie group, the rank of a closed subgroup cannot exceed the rank of the ambient group. Since $\text{rank}(SU(n))=n-1$, an embedding of the SM gauge group into $SU(n)$ requires $n-1\geq 4$, hence $n\geq 5$.
The choice $n=5$ is, therefore, the minimal rank allowing an $SU(n)$ internal structure compatible with an $SU(3)\times SU(2)\times U(1)$ subgroup. The standard $SU(5)$ embedding is realized by an adjoint breaking with hypercharge generator:
$Y\propto\mathrm{diag}\!\left(-\tfrac{1}{3},-\tfrac{1}{3},-\tfrac{1}{3},\tfrac{1}{2},\tfrac{1}{2}\right),$ (14)
and chirality or anomaly cancellation for one family is then achieved by the usual $\mathbf{10}\oplus\overline{\mathbf{5}}$ matter assignment with the additional geometric integrality condition fixing the $U(1)$ charge lattice. We can go over $\text{SU}(5)$ from geometry in greater detail in a future paper. Therefore, the maximal compatible internal symmetry is
@@
under a holomorphic adjoint ($\mathbf{24}$) reduction, with the standard hypercharge generator GeorgiGlashow1974 ; Slansky1981 . In what follows, we take $G=\mathrm{SU}(5)$ above the mass-energy scale $M_{*}$, with a holomorphic adjoint reduction to the Standard Model group on the real slice. Given the rank-$n$ holomorphic bundle $E\to M$ with the Hermitian metric on the internal vector bundle $h_{E}$ and a nowhere-vanishing unitary volume form $\Omega_{E}\in H^{0}(M,\wedge^{n}E^{*})$, the structure group reduces as follows:
@@
The preservation of $h_{E}$ picks unitary changes of frame, such as $\mathrm{U}(n)\subset\mathrm{GL}(n,\mathbb{C})$. Preservation of $\Omega_{E}$ forces $\det=1$ on $\mathrm{U}(n)$, giving $\mathrm{SU}(n)$. If, in addition, $c_{1}(E)=0$ and $E$ is slope-stable, the unique compatible connection is Hermitian–Yang–Mills and its holonomy lies in $\mathrm{SU}(n)$ Donaldson1985 ; UhlenbeckYau1986 . The internal gauge symmetry is the automorphism group preserving $(h_{E},\Omega_{E})$, namely $G=\mathrm{SU}(n)$. Imposing chirality and anomaly cancellation for one SM family, hypercharge integrality via $\frac{1}{8\pi^{2}}\int\mathrm{Tr}(F_{Y}\wedge F_{Y})\in\mathbb{Z}$, and a single adjoint breaking on the real slice fixes the minimal rank to $n=5$, we have: